# Physical Correlations and Predictions Emerging from Modern Core-Collapse Supernova Theory

### 12 Jan 2024 v1

@@ 5.4 Kicks
Burrows et al. (2023b) have recently crafted and submitted a paper on the physics and systematics of the recoil kicks and induced spins, and their potential correlations, for our long-term 3D simulations. Though that paper contains the same models, the time between this paper and that has allowed us to carry those models even further. However, the general conclusions in Burrows et al. (2023b) are unaltered and we won’t repeat its discussion and conclusions in detail here. We found an important correlation between the kick speed and both the energy of explosion and the ejecta mass dipole. We identified a weakly inverse relation between induced spin and kick speed, with a lot of scatter, and found a weak induced-spin/kick correlation. If the initial rotational vector of the collapsing core correlates somewhat with the direction of explosion, a spin/kick correlation would naturally emerge (Holland-Ashford et al., 2017; Katsuda et al., 2018; Ng & Romani, 2007). We also found a compelling explanation for the lower average gravitational masses of binary neutron stars in the neutron star mass/kick connection that emerges from the simulations, viewed collectively. The lowest mass progenitors give birth to the lowest mass neutron stars with the slowest kicks, and this would make binary disruption harder.
Therefore, in this paper we add only two more plots that reveal two additional correlations. Figure 17 indicates that the kick speed and the energy dipole of the ejecta (left) and the combination (right) (where E is the explosion energy and is the baryon mass of the neutron star) are tightly correlated. The more asymmetrical the explosion the greater the kick speed. Moreover, the kick speed is related to the explosion energy itself, divided by the residual neutron star mass. In Burrows et al. (2023b), we found similar relations between kick speed and explosion energy and ejecta mass dipole. However, the correlation between kick speed and ejecta energy dipole is actually a bit tighter. In any case, the connections between kick speed and various measures of anisotropy, while not unexpected, are testable.
@@ 6 Conclusions
In this paper, we have derived correlations between core-collapse supernova observables and progenitor core structures that naturally emerge from our extensive suite of twenty state-of-the-art 3D CCSN simulations carried to late times. This is the largest such collection of 3D supernova models ever generated and allows one to witness and derive patterns that might otherwise be obscured when studying one or a few models in isolation. Moreover, such long-term simulations are necessary to capture the asymptotic values of many of the most important observable quantities, and such simulations have until recently been too expensive to generate. However, with the efficiency of our new supernova code Fornax and recent generous allocations of supercomputer time, this limitation has now been overcome. Furthermore, a large collection of such models allows one partially to mitigate against the fact that the physical turbulence and chaos that attends supernova dynamics will naturally translate into a distribution of outcomes and explosion parameters in deriving trends.555The associated scatter and spread in observables is unknown and is properly a subject of future research.. Looking at single models one at a time tends to obscure the overall correlations, with both progenitor structure and between observables.
@@
From this new panoramic perspective, we have discovered 1) observationally useful correlations between explosion energy, neutron star gravitational birth masses, Ni and -rich freeze-out yields, and pulsar kicks (Burrows et al., 2023b) and 2) physically and theoretically important correlations with the compactness parameter of progenitor structure. The theoretical correlation between Ni and explosion energy has been seen before (e.g. Sukhbold et al. (2016)), but not in the context of such a large 3D model suite carried to, or near to, the asymptotic state. Such a correlation is also strongly suggested by observations (Rodríguez et al., 2023). The intriguing correlation between explosion energy and neutron star gravitational mass is new. In addition, we find a correlation between explosion energy and progenitor mantle binding energy, suggesting that neutrino-driven explosions are self-regulating. We also find a testable correlation between explosion energy and measures of explosion asymmetry, such as the ejecta energy and mass dipoles.
@@
While the correlations between two observables are roughly independent of the progenitor ZAMS mass, the many correlations we have derived with compactness can not unambiguously be tied to a particular progenitor ZAMS mass. The compactness/ZAMS mass mapping depends upon the massive star progenitor models employed and the mapping between compactness and/or initial density structure and progenitor mass is still in flux. Therefore, our derived correlations between compactness and observables may be more robust than with ZAMS mass. There remain may important issues, such as binarity, shell mixing, overshoot, wind mass loss, semi-convection, the C(,) nuclear rate, and Roche-lobe overflow, that remain to be resolved related to the endpoint structures of massive stars. Hence, it is likely that the mapping between compactness and ZAMS mass found in the Sukhbold et al. (2016) and Sukhbold et al. (2018) compendia is not the final word.
@@
Though Fornax is a sophisticated, multi-physics multi-group radiation/hydrodynamics code that has been matured over the last eight years to address the complicated multi-dimensional phenomena encountered in core collapse, there are numerous upgrades and improvements that suggest themselves. We are bundling the , , neutrinos into one pseudo-species. This is done because their neutrino-matter interactions are quite similar and to trim the computational load. However, their interactions are not exactly the same (Bollig et al., 2017). Also, we are using 12 (3) energy groups and though we have in the past determined that this number is adequate, more would be better. In the same spirit, an augmentation of the number of zones and enhanced spatial resolution is always good, if affordable. Currently, we are using a 1024()128()256() grid and this is better than most 3D production simulations. However, we have shown that resolution can make a difference, perhaps qualitatively (Nagakura et al., 2019). Importantly, we are using approximate general relativity (Marek et al., 2006), which captures the basic differences in the strength of gravity between the Newtonian and Einsteinian realizations, provides an accurate lapse function, and allows for the incorporation of gravitational redshifts. However, full general relativity is the correct approach. Also, we have not attempted to capture the effects of neutrino oscillations and we are wedded in this simulation suite to the Sukhbold et al. (2016) and Sukhbold et al. (2018) initial models. The latter are certainly not the last word, are not 3D, and do not include the many effects of binarity. Furthermore, as we have demonstrated, the initial structures are crucial to determining the mapping between star and outcome and we have not incorporated the effects of magnetic fields (Müller & Varma, 2020; Varma & Müller, 2021; Obergaulinger & Aloy, 2020; Kuroda et al., 2020; Obergaulinger & Aloy, 2021), nor addressed the effects of initial core rotation. Finally, the chaos of the turbulent context of post-bounce evolution introduces a degree of randomness and stochasticity that will translate into distribution functions for the observables, even for the same progenitor, whose widths are currently unknown. Hence, one can envision many potential upgrades to current computational platforms, and many necessary and more detailed investigations to pursue, in order to further improve the 3D theory of core-collapse supernova explosions. Nevertheless, our current collection of twenty state-of-the-art 3D models provides a synoptic view unavailable in the past upon which future work can profitably build.
@@ Acknowledgments
Fornax is written in C. We use an MPI programming model for the CPU variant and an MPI/OpenMP-4.5/5.0 hybrid paralellism model for the GPU variant. The code employs spherical coordinates in one, two, and three spatial dimensions, solves the comoving-frame, multi-group, two-moment, velocity-dependent transport equations, and uses the M1 tensor closure for the second and third moments of the radiation fields Vaytet et al. (2011)666Most, though not all, supernova groups in the world employ the so-called “ray-by-ray+” dimensional reduction approach to transport (Bruenn et al., 2020; Melson et al., 2020). This is not true multi-D transport, but multiple spherical (1D) solves, and is done for speed. However, the results for multi-D evolutions that are not spherical can be in qualitative error (Skinner et al., 2016).. Three species of neutrino (, , and “” [, , , and lumped together]) are followed using an explicit Godunov characteristic method applied to the radiation transport operators, but an implicit solver for the radiation source terms. This solver for the transport operators can be explicit since the speed of light and the speed of sound in the inner core are close and we are already timestep-limited by the hydro Courant condition (Just et al., 2015; O’Connor, 2015; Skinner et al., 2019). By addressing the transport operator with an explicit method, we significantly reduce the computational complexity and communication overhead of traditional multi-dimensional radiative transfer solutions by bypassing the need for global iterative solvers. Radiation quantities are reconstructed with linear profiles and the calculated edge states are used to determine fluxes via an HLLE solver. In the non-hyperbolic regime, the HLLE fluxes are corrected to reduce numerical diffusion (O’Connor & Ott, 2013). The momentum and energy transfer between the radiation and the gas are operator-split and addressed implicitly.

### 12 Mar 2024 v2

@@ 5.4 Kicks
Burrows et al. (2023b) have recently crafted and had accepted a paper on the physics and systematics of the recoil kicks and induced spins, and their potential correlations, for our long-term 3D simulations. Though that paper contains the same models, the time between this paper and that has allowed us to carry those models even further. However, the general conclusions in Burrows et al. (2023b) are unaltered and we won’t repeat its discussion and conclusions in detail here. We found an important correlation between the kick speed and both the energy of explosion and the ejecta mass dipole. We identified a weakly inverse relation between induced spin and kick speed, with a lot of scatter, and found a weak induced-spin/kick correlation. If the initial rotational vector of the collapsing core correlates somewhat with the direction of explosion, a spin/kick correlation would naturally emerge (Holland-Ashford et al., 2017; Katsuda et al., 2018; Ng & Romani, 2007). We also found a compelling explanation for the lower average gravitational masses of binary neutron stars in the neutron star mass/kick connection that emerges from the simulations, viewed collectively. The lowest mass progenitors give birth to the lowest mass neutron stars with the slowest kicks, and this would make binary disruption harder.
Therefore, in this paper we add only two more plots that reveal two additional correlations. Figure 17 indicates that the kick speed and the energy dipole of the ejecta (left) and the combination (right) (where E is the explosion energy and is the baryon mass of the neutron star) are tightly correlated. The more asymmetrical the explosion the greater the kick speed. Moreover, the kick speed is related to the explosion energy itself, divided by the residual neutron star mass. In Burrows et al. (2023b), we found similar relations between kick speed and explosion energy and ejecta mass dipole, as did Janka (2017b)666However, we note that Janka (2017b) “calibrates” his alpha parameter on the Scheck et al. (2006) and Wongwathanarat et al. (2013) papers. The former is 2D and the latter used a light bulb to artificially explode models. Moreover, Wongwathanarat et al. (2013) found a “reduced ejecta asymmetry with higher explosion energies,” something we do not see.. However, the correlation between kick speed and ejecta energy dipole is actually a bit tighter. In any case, the connections between kick speed and various measures of anisotropy, while not unexpected, are testable.
@@ 6 Conclusions
In this paper, we have derived correlations between core-collapse supernova observables and progenitor core structures that naturally emerge from our extensive suite of twenty state-of-the-art 3D CCSN simulations carried to late times. This is the largest such collection of 3D supernova models ever generated and allows one to witness and derive patterns that might otherwise be obscured when studying one or a few models in isolation. Moreover, such long-term simulations are necessary to capture the asymptotic values of many of the most important observable quantities, and such simulations have until recently been too expensive to generate. However, with the efficiency of our new supernova code Fornax and recent generous allocations of supercomputer time, this limitation has now been overcome. Furthermore, a large collection of such models allows one partially to mitigate against the fact that the physical turbulence and chaos that attends supernova dynamics will naturally translate into a distribution of outcomes and explosion parameters in deriving trends.777The associated scatter and spread in observables is unknown and is properly a subject of future research.. Looking at single models one at a time tends to obscure the overall correlations, with both progenitor structure and between observables.
@@
From this new panoramic perspective, we have identified 1) observationally useful correlations between explosion energy, neutron star gravitational birth masses, Ni and -rich freeze-out yields, and pulsar kicks (Burrows et al., 2023b) and 2) physically and theoretically important correlations with the compactness parameter of progenitor structure. The theoretical correlation between Ni and explosion energy has been seen before (e.g. Sukhbold et al. (2016)), but not in the context of such a large 3D model suite carried to, or near to, the asymptotic state. Such a correlation is also strongly suggested by observations (Rodríguez et al., 2023). The intriguing correlation between explosion energy and neutron star gravitational mass is new in the detailed 3D modeling context, but was anticipated in the 2D results of Nakamura et al. (2015) and in the semi-analytic modeling of Müller et al. (2016a). In addition, we find a correlation between explosion energy and progenitor mantle binding energy, suggesting that neutrino-driven explosions are self-regulating. We also find a testable correlation between explosion energy and measures of explosion asymmetry, such as the ejecta energy and mass dipoles.
@@
While the correlations between two observables are roughly independent of the progenitor ZAMS mass, the many correlations we have derived with compactness can not unambiguously be tied to a particular progenitor ZAMS mass. The compactness/ZAMS mass mapping depends upon the massive star progenitor models employed and the mapping between compactness and/or initial density structure and progenitor mass is still in flux. Therefore, our derived correlations between compactness and observables may be more robust than with ZAMS mass. There remain may important issues, such as binarity, shell mixing, overshoot, wind mass loss, semi-convection, the C(,) nuclear rate, and Roche-lobe overflow, that remain to be resolved related to the endpoint structures of massive stars. Hence, it is likely that the mapping between compactness and ZAMS mass found in the Sukhbold et al. (2016) and Sukhbold et al. (2018) compendia is not the final word. Using a simple analytic model for core-collapse supernovae and estimates of some explosion observables, this point is clearly made in Temaj et al. (2023).
@@
Though Fornax is a sophisticated, multi-physics multi-group radiation/hydrodynamics code that has been matured over the last eight years to address the complicated multi-dimensional phenomena encountered in core collapse, there are numerous upgrades and improvements that suggest themselves. We are bundling the , , neutrinos into one pseudo-species. This is done because their neutrino-matter interactions are quite similar and to trim the computational load. However, their interactions are not exactly the same (Bollig et al., 2017). Also, we are using 12 (3) energy groups and though we have in the past determined that this number is adequate, more would be better. In the same spirit, an augmentation of the number of zones and enhanced spatial resolution are always good, if affordable. Currently, we are using a 1024()128()256() grid and this is better than most 3D production simulations. However, we have shown that resolution can make a difference, perhaps qualitatively (Nagakura et al., 2019). Importantly, we are using approximate general relativity (Marek et al., 2006), which captures the basic differences in the strength of gravity between the Newtonian and Einsteinian realizations, provides an accurate lapse function, and allows for the incorporation of gravitational redshifts. However, full general relativity is the correct approach. Also, we have not attempted to capture the effects of neutrino oscillations and we are wedded in this simulation suite to the Sukhbold et al. (2016) and Sukhbold et al. (2018) initial models. The latter are certainly not the last word, are not 3D, and do not include the many effects of binarity. Furthermore, as we have demonstrated, the initial structures are crucial to determining the mapping between star and outcome and we have not incorporated the effects of magnetic fields (Müller & Varma, 2020; Varma & Müller, 2021; Obergaulinger & Aloy, 2020; Kuroda et al., 2020; Obergaulinger & Aloy, 2021), nor addressed the effects of initial core rotation. Finally, the chaos of the turbulent context of post-bounce evolution introduces a degree of randomness and stochasticity that will translate into distribution functions for the observables, even for the same progenitor, whose widths are currently unknown. Hence, one can envision many potential upgrades to current computational platforms, and many necessary and more detailed investigations to pursue, in order to further improve the 3D theory of core-collapse supernova explosions. Nevertheless, our current collection of twenty state-of-the-art 3D models provides a synoptic view unavailable in the past upon which future work can profitably build.
@@ Acknowledgments
Fornax is written in C. We use an MPI programming model for the CPU variant and an MPI/OpenMP-4.5/5.0 hybrid paralellism model for the GPU variant. The code employs spherical coordinates in one, two, and three spatial dimensions, solves the comoving-frame, multi-group, two-moment, velocity-dependent transport equations, and uses the M1 tensor closure for the second and third moments of the radiation fields Vaytet et al. (2011)888Most, though not all, supernova groups in the world employ the so-called “ray-by-ray+” dimensional reduction approach to transport (Bruenn et al., 2020; Melson et al., 2020). This is not true multi-D transport, but multiple spherical (1D) solves, and is done for speed. However, the results for multi-D evolutions that are not spherical can be in qualitative error (Skinner et al., 2016, in 2D). Glas et al. (2019) explored the same issues in 3D and showed that when the explosion is roughly spherical (as in their 9 model) the ray-by-ray+ approach is adequate. However, they make a less convincing argument when the explosion is aspherical (as in their 20 model).. Three species of neutrino (, , and “” [, , , and lumped together]) are followed using an explicit Godunov characteristic method applied to the radiation transport operators, but an implicit solver for the radiation source terms. This solver for the transport operators can be explicit since the speed of light and the speed of sound in the inner core are close and we are already timestep-limited by the hydro Courant condition (Just et al., 2015; O’Connor, 2015; Skinner et al., 2019). By addressing the transport operator with an explicit method, we significantly reduce the computational complexity and communication overhead of traditional multi-dimensional radiative transfer solutions by bypassing the need for global iterative solvers. Radiation quantities are reconstructed with linear profiles and the calculated edge states are used to determine fluxes via an HLLE solver. In the non-hyperbolic regime, the HLLE fluxes are corrected to reduce numerical diffusion (O’Connor & Ott, 2013). The momentum and energy transfer between the radiation and the gas are operator-split and addressed implicitly.